# Feature-based Low-Rank Compression of Large Language Models via Bayesian Optimization

### v1

@@ Abstract
Feature-based Low-Rank Compression of Large Language Models via Bayesian Optimization
Yixin Ji1thanks:   Equal Contribution., Yang Xiang [^1] footnotemark: 1, Juntao Li1thanks:   Corresponding author., Wei Chen2, Zhongyi Liu2, Kehai Chen3, Min Zhang1 1School of Computer Science and Technology, Soochow University 2Ant Group 3Harbin Institute of Technology, Shenzhen {jiyixin169,baldwin021129}@gmail.com; {qianmu.cw, chonghui.lzy}@antgroup.com; chenkehai@hit.edu.cn; {ljt,minzhang}@suda.edu.cn
@@ 2.1 Weight-based and Feature-based Low-rank Decomposition
The low-rank decomposition reduces the number of parameters by decomposing the linear layer weights into two low-rank matrices. Weight-based factorization is one naive method. For a linear layer $\boldsymbol{W}\in\mathbb{R}^{d_{2}\times d_{1}}$, the objective can be formulated as:
$\underset{\boldsymbol{B},\boldsymbol{A}}{\text{min}}||\boldsymbol{W}-\boldsymbol{B}\boldsymbol{A}||_{F}$ (1)
[…]
If $r<d_{1}d_{2}/(d_{1}+d_{2})$, the factorization can reduce the total parameter amount. According to the Eckart–Young–Mirsky theorem, the trunced singular value decomposition (SVD) provides the optimal solution:
$\boldsymbol{W}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{T}$ (2)
$\boldsymbol{A}=\boldsymbol{V}_{r}^{T},\boldsymbol{B}=\boldsymbol{U}_{r}\boldsymbol{\Sigma}_{r},$
where $\boldsymbol{A}\in\mathbb{R}^{r\times d_{1}},\boldsymbol{B}\in\mathbb{R}^{d_{2}\times r}$, $\boldsymbol{\Sigma}_{r}$ is the top-$r$ largest singular values, $\boldsymbol{U}_{r}$ and $\boldsymbol{V}_{r}$ are the corresponding singular vectors.
However, in the vast majority of cases, the weights of PLMs have a high rank, and a direct truncated SVD decomposition on the weights would lead to significant reconstruction errors Chen et al. (2021). In comparison, the representation space of PLMs exhibits a clear low-rank property Yu and Wu (2023). Therefore, another line of work has considered feature-based factorization:
@@ 2.2 Different Layers Exhibit Varying Degrees of Low-rank Sensitivity
Another challenge in LRC is allocating varying low-rank compression rates to different layers. Previous works have empirically or theoretically demonstrated that different components of Transformer-based masked language models and visual models exhibit distinct low-rank properties, such as the features of the self-attention modules having a lower rank compared to those of the feed-forward modules Dong et al. (2023); Anagnostidis et al. (2022). These findings provide prior guidance for low-rank compression. However, detailed studies on current mainstream large language models are still lacking. Therefore, we take the LLaMA-v2-7b as an example to study the low-rank sensitivity within each layer across different types of layers and the same type of layers. Llama-family LLMs have seven distinct parameter categories: attn_q, attn_k, attn_v, attn_o, mlp_up, mlp_down, and mlp_gate. We evaluate the perplexity changes on Wikitext-2 Merity et al. (2016) for each category under varying low-rank compression ratios. As Figure 1 shows, at the same low-rank compression rate, distinct layers exhibit notable performance variations. For attn_q and attn_k, they demonstrate robustness to low-rank compression, with an increase in perplexity not exceeding 2% even at a compression rate of 60%. In contrast, attn_v, with an equivalent parameter count, exhibits high sensitivity, leading to a significant surge in perplexity with compression rates even below 5%. Therefore, assigning the same low-rank compression rate to different types of layers during low-rank compression of LLM is a sub-optimal solution. More empirical study results are shown in Appendix A.
@@ 3.2 Low-Rank Allocation Based on Bayesian Optimization
As investigated in Section 2.2, different types of layers, and even each individual layer, exhibit varying sensitivities to low-rank compression. Therefore, allocating distinct compression ratios to different layers is crucial to achieve the desired compression rate with minimal performance degradation. For a LLM $f(\cdot;\boldsymbol{\theta})$, we compress it with the set of low-rank compression ratios $\boldsymbol{\lambda}=\{\lambda_{i}\}_{i=1}^{k}$. We use a task-agnostic evaluation dataset $\mathcal{D}$ to evaluate performance of the compressed model $f(\cdot;\boldsymbol{\theta},\boldsymbol{\lambda})$. For example, we can take the perplexity on a subset of pretraining data as the evaluation metric. Therefore, the optimization objective of low-rank allocation can be formulated as:
@@
where $h(\cdot,\cdot)$ is the evaluation metric, $\rho$ is model’s overall compression ratio. For LLMs, finding the optimal low-rank allocation based on a target compression ratio is a challenging optimization problem. First, the impact of the low-rank count allocated to different layers on the performance of the compressed model is combinatorial, and optimizing any one component independently may lead to a locally optimal solution. Second, due to LLMs’ vast number of parameters, evaluating $H(\boldsymbol{\lambda})$ is very time-consuming. Therefore, we leverage sample-efficient Bayesian optimization Xu et al. (2022) to optimize Eq 7. Bayesian optimization (BO) estimates the objective $H(\boldsymbol{\lambda})$ using a stochastic surrogate model and updates the posterior estimation of $H(\boldsymbol{\lambda})$ based on the results of each search step. We utilize the Gaussian process $\mathcal{N}(\mu(\cdot),\sigma^{2}(\cdot))$ as the surrogate model. Given the previous $t-1$ search steps $\{\boldsymbol{\lambda}_{1},\cdots,\boldsymbol{\lambda}_{t-1}\}$ and their evaluation $H_{t-1}=[H(\boldsymbol{\lambda}_{1}),\cdots,H(\boldsymbol{\lambda}_{t-1})]$, the surrogate model is updated as:
@@
where $H^{\prime}=\min_{i\in[t-1]}H(\boldsymbol{\lambda}_{i})$, it means the minimal value observed so far. Then, BO chooses the point with the greatest expected improvement to explore. After obtaining the optimal ratio $\boldsymbol{\lambda}^{*}$, we can determine the low rank. To fully leverage the acceleration effect of GPU matrix multiplication, we adhere to Nvidia’s user guidelines by rounding the low-rank dimensions to the nearest multiple of eight [^2] https://docs.nvidia.com/deeplearning/performance/dl-performance-matrix-multiplication/index.html#gpu-imple:
$r_{i}=[\frac{(1-\lambda_{i})d_{1}d_{2}}{d_{1}+d_{2}}/8]*8.$ (10)
During the Bayesian optimization process, to minimize the evaluation cost, we use a smaller validation set that may not comprehensively reflect the LLM’s performance, potentially leading to over-fitting in the validation set during the optimization. To prevent BO from blindly improving the compressed model’s language modeling performance on the validation set, we aim to make the compressed model to have a prediction distribution for the next word that is close to that of the original model. Hence, we employ the reverse KL divergence to quantify the difference:
@@ 4 Experiments
4.1 Baseline Methods
We compare our method with the following competitive structured pruning and low-rank compression methods in LLMs.
LLM-Pruner Ma et al. (2023) is a dependency-aware one-shot structured pruning. It evaluates the importance of each structure through a first-order Taylor expansion and prunes the structures with the lowest scores. After pruning, it uses LoRA post-training to recover performance.
4.2 Datasets
@@ 4.3 Experimental Details
In our main experiments, we apply our method to LLaMA-v2-7b and LLaMA-v2-13b. We randomly select 1,024 samples from the training set of C4 as the calibration data. Each sample has a sequence length of 4,096. To estimate the covariance matrix while saving memory usage, we employ the Welford’s online algorithm Welford (1962). For the pooled covariance matrix, we partition the calibrated data into 32 groups. During the Bayesian optimization, we utilize the Matern kernel as the covariance function and use the test set of WikiText2 (about 34k tokens) as the evaluation data. Considering that Bayesian optimization is not well-suited for high-dimensional scenarios, we conduct experiments with two settings based on the observations in Section 2.2: (a) $5\times 1$: We allow $attn\_q$ and $attn\_k$ to share a low-rank dimension, and the same type of parameters across different layers to also share a low-rank dimension, thus BO only optimizes 5 parameters; (b) $5\times 4$: Building on the setup of (a), we divide the model’s layers into 4 groups in sequence, with no parameter sharing between different groups, resulting in BO needing to optimize 20 parameters. We run 50 epochs of Bayesian optimization to search the optimal low-rank allocation. At the post-training stage, following LLM-Pruner, we use the Alpaca dataset Taori et al. (2023) and train 2 epochs. More details can be found in the Appendix B.
@@ 4.4 Main Results
Bolaco ($5\times 1$) 71.44 75.24 67.22 67.09 72.85 39.16 42.00 62.14
Bolaco ($5\times 1$ w/ PT) 74.56 76.61 71.74 66.77 73.57 42.75 43.20 64.17
Bolaco ($5\times 4$) 71.44 74.70 67.24 66.54 71.84 38.31 42.60 61.81
Bolaco ($5\times 4$ w/ PT) 75.90 76.44 71.85 67.09 72.98 41.89 42.20 64.05
@@
Bolaco ($5\times 1$) 69.11 71.55 56.02 63.69 64.52 33.02 38.40 56.62
Bolaco ($5\times 1$ w/ PT) 70.80 74.81 67.38 64.64 68.35 38.65 41.00 60.80
Bolaco ($5\times 4$) 73.85 71.27 57.85 64.48 65.87 34.30 37.80 57.92
Bolaco ($5\times 4$ w/ PT) 71.01 74.59 67.68 65.75 69.07 38.99 40.80 61.13
0% LLaMA-v2-13b 80.52 79.05 79.38 72.14 79.42 49.23 45.20 69.27
20% LLM-Pruner 66.33 78.18 74.47 64.48 72.26 45.90 44.20 63.69
LLM-Pruner (w/ PT) 67.06 78.94 75.92 67.32 72.69 44.28 44.60 64.40
FLAP 71.28 76.55 74.67 69.53 72.56 44.03 42.00 64.37
LoRD 78.47 76.01 69.58 71.03 74.33 40.87 44.40 64.96
Bolaco ($5\times 1$) 80.58 77.04 73.11 70.01 76.39 44.28 45.00 66.63
Bolaco ($5\times 1$ w/ PT) 81.25 78.51 76.70 71.35 77.10 47.44 44.80 68.16
Bolaco ($5\times 4$) 78.81 76.33 72.50 69.53 75.42 42.23 45.00 65.69
Bolaco ($5\times 4$ w/ PT) 79.33 78.07 76.41 70.80 75.29 44.80 44.80 67.07
30% LLM-Pruner 62.45 75.90 67.90 60.22 65.45 40.36 44.60 59.55
LLM-Pruner (w/ PT) 68.29 76.66 72.03 64.09 69.20 41.13 45.40 62.40
FLAP 65.54 74.81 70.29 67.48 67.38 38.23 40.00 60.53
LoRD 75.05 73.88 63.08 69.46 69.78 39.16 38.60 61.29
Bolaco ($5\times 1$) 76.42 74.86 64.36 67.01 71.93 40.19 40.60 62.20
Bolaco ($5\times 1$ w/ PT) 79.39 76.33 73.21 69.46 74.12 43.77 44.20 65.78
Bolaco ($5\times 4$) 75.63 74.16 66.66 69.06 71.76 40.10 41.60 62.71
Bolaco ($5\times 4$ w/ PT) 80.34 75.79 73.98 67.88 73.40 45.14 44.60 65.88
Table 1: Zero-shot performance of the compressed LLaMA-v2 models. w/ PT means the method with post-training. Bold denotes the best result at the same compression ratio, while underline indicates the second best result.
We report the perplexity of language modeling for various compression methods at different compression ratios in Figure 3 and 6, and the zero-shot common sense reasoning results in Table 1. In terms of language modeling capabilities, FLAP demonstrates strong competitiveness, particularly when the compression rate exceeds 30%, where FLAP’s perplexity is slightly better than our Bolaco ($5\times 1$). However, in the 7b model, Bolaco ($5\times 4$) achieves the best language modeling performance at high compression rates. Nevertheless, in the 13b model, despite Bolaco ($5\times 4$) still leading other compression techniques, it maintains a certain gap from FLAP. For zero-shot tasks, our method significantly outperforms all baselines without any further post-training, achieving an average performance increase of 1.5-2% across seven datasets. After post-training with only about 1‰ parameters and 3 hours, our method further narrows down the performance difference between the compressed model and the original model. It retains 96%-98% of the original model’s performance at the 20% compression ratio, and at a 30% compression ratio, it maintains 91%-95% of the performance. Comparing the $5\times 1$ and $5\times 4$ setting, we find that the performance difference between the two is not significant. At the 20% compression ratio, simply allocating different low-rank dimensions to different types of parameters suffices to achieve the best current performance. However, at the 30% compression rate, the $5\times 4$ setting outperforms the $5\times 1$, indicating that more granular low-rank assignments contribute to enhanced performance in compressed models at higher compression rates.
@@ 5.2 Impact of Objective Function
PPL ($5\times 1$) 8.04 48.94 61.83
w/ RKL ($5\times 1$) 7.61 48.37 62.14
PPL ($5\times 4$) 7.78 48.37 61.05
w/ RKL ($5\times 4$) 8.08 48.85 61.81
@@
PPL ($5\times 1$) 13.80 76.70 56.84
w/ RKL ($5\times 1$) 13.91 81.64 56.62
PPL ($5\times 4$) 12.74 77.30 56.32
w/ RKL ($5\times 4$) 11.99 71.49 57.92
@@
We explore the impact of the objective function in the Bayesian optimization stage. We conduct experiments on LLaMA-v2-7b and report results in Table 3. Overall, incorporating the reverse KL divergence (RKL) between the compressed model and the original model’s predictive distribution into the objective function can lead to a better low-rank dimensions allocation. Especially in the $5\times 4$ setting, which is more difficult to optimize for Bayesian optimization, the performance gains from RKL term are even more obvious. We suppose that the RKL term may serve two roles. Firstly, as a regularization term, it prevents overfitting on smaller validation sets during Bayesian optimization. Although the compressed model exhibits a slight increase in perplexity on the language modeling dataset at the 20% compression rate with the $5\times 4$ setting, there is a significant improvement in performance on downstream tasks. Secondly, incorporating the RKL term may smooth the objective function, enabling the Gaussian process surrogate model to more accurately approximate the real black-box objective function.

### v2

# Adaptive Feature-based Low-Rank Compression of Large Language Models via Bayesian Optimization

@@ Abstract
Adaptive Feature-based Low-Rank Compression of Large Language Models via Bayesian Optimization
Yixin Ji1thanks:   Equal Contribution., Yang Xiang [^1] footnotemark: 1, Juntao Li1thanks:   Corresponding author., Qingrong Xia2, Zi Ye2, Xinyu Duan2, Zhefeng Wang2, Kehai Chen3, Min Zhang1,3 1School of Computer Science and Technology, Soochow University 2Huawei Cloud, China 3Harbin Institute of Technology, Shenzhen {jiyixin169,baldwin021129}@gmail.com; {ljt,minzhang}@suda.edu.cn
@@ 2.1 Weight-based and Feature-based Low-rank Decomposition
The low-rank decomposition reduces the number of parameters by decomposing the linear layer weights into two low-rank matrices. Weight-based factorization is one naive method. For a linear layer $\boldsymbol{W}\in\mathbb{R}^{d_{2}\times d_{1}}$, according to the Eckart–Young–Mirsky theorem, the trunced singular value decomposition (SVD) provides the optimal solution: $\boldsymbol{W}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{T},\boldsymbol{A}=\boldsymbol{V}_{r}^{T},\boldsymbol{B}=\boldsymbol{U}_{r}\boldsymbol{\Sigma}_{r},$ where $\boldsymbol{A}\in\mathbb{R}^{r\times d_{1}},\boldsymbol{B}\in\mathbb{R}^{d_{2}\times r}$, $\boldsymbol{\Sigma}_{r}$ is the top-$r$ largest singular values, $\boldsymbol{U}_{r}$ and $\boldsymbol{V}_{r}$ are the corresponding singular vectors. If $r<d_{1}d_{2}/(d_{1}+d_{2})$, the factorization can reduce the total parameter amount. However, in the vast majority of cases, the weights of PLMs have a high rank, and a direct truncated SVD decomposition on the weights would lead to significant reconstruction errors Chen et al. (2021). In comparison, the representation space of PLMs exhibits a clear low-rank property Yu and Wu (2023). Therefore, another line of work has considered feature-based factorization:
@@ 2.2 Different Layers Exhibit Varying Degrees of Low-rank Sensitivity
Another challenge in LRC is allocating varying low-rank compression rates to different layers. Previous works have empirically or theoretically demonstrated that different components of Transformer-based masked language models and visual models exhibit distinct low-rank properties, such as the features of the self-attention modules having a lower rank compared to those of the feed-forward modules Dong et al. (2023); Anagnostidis et al. (2022). These findings provide prior guidance for low-rank compression. However, detailed studies on current mainstream LLMs are still lacking. Therefore, we take the LLaMA-v2-7b as an example to study the low-rank sensitivity within each layer across different types of layers and the same type of layers. Llama-family LLMs have seven distinct parameter categories: attn_q, attn_k, attn_v, attn_o, mlp_up, mlp_down, and mlp_gate. We evaluate the perplexity changes on Wikitext-2 Merity et al. (2016) for each category under varying low-rank compression ratios. As Figure 1 shows, at the same low-rank compression rate, distinct layers exhibit notable performance variations. For attn_q and attn_k, they demonstrate robustness to low-rank compression, with an increase in perplexity not exceeding 2% even at a compression rate of 60%. In contrast, attn_v, with an equivalent parameter count, exhibits high sensitivity, leading to a significant surge in perplexity with compression rates even below 5%. Therefore, assigning the same low-rank compression rate to different types of layers during low-rank compression of LLM is a sub-optimal solution. In addition to the differences, we also observe certain similarities, e.g., attn_q and attn_k have similar low-rank sensitivities. More empirical study results are shown in Appendix A.
@@ 3.2 Low-Rank Allocation Based on Bayesian Optimization
As investigated in Section 2.2, different types of layers, and even each individual layer, exhibit varying sensitivities to low-rank compression. Therefore, allocating distinct compression ratios to different layers is crucial to achieve the desired compression rate with minimal performance degradation. For a LLM $f(\cdot;\boldsymbol{\theta})$, we compress it with the set of low-rank compression ratios $\boldsymbol{\lambda}=\{\lambda_{i}\}_{i=1}^{k}$. We use a task-agnostic evaluation dataset $\mathcal{D}$ to evaluate performance of the compressed model $f(\cdot;\boldsymbol{\theta},\boldsymbol{\lambda})$, such as the perplexity on a subset of pretraining data. Therefore, the optimization objective of low-rank allocation can be formulated as:
@@
where $h(\cdot,\cdot)$ is the evaluation metric, $\rho$ is model’s overall compression ratio. For LLMs, searching the optimal low-rank allocation is a challenging optimization problem. First, the impact of the low-rank count allocated to different layers on the performance of the compressed model is combinatorial, and optimizing any one component independently may lead to a locally optimal solution. Then, due to LLMs’ vast number of parameters, evaluating $H(\boldsymbol{\lambda})$ is very time-consuming. Therefore, we leverage sample-efficient Bayesian optimization (BO) Xu et al. (2022) to optimize Eq 5. BO estimates the objective $H(\boldsymbol{\lambda})$ with a stochastic surrogate model and updates the posterior estimation of $H(\boldsymbol{\lambda})$ based on the results of each search step. We utilize the Gaussian process $\mathcal{N}(\mu(\cdot),\sigma^{2}(\cdot))$ as the surrogate model. Given the previous $t-1$ search steps $\{\boldsymbol{\lambda}_{1},\cdots,\boldsymbol{\lambda}_{t-1}\}$ and their evaluation $H_{t-1}=[H(\boldsymbol{\lambda}_{1}),\cdots,H(\boldsymbol{\lambda}_{t-1})]$, the surrogate model is updated as:
@@
where $H^{\prime}=\min_{i\in[t-1]}H(\boldsymbol{\lambda}_{i})$, it means the minimal value observed so far. Then, BO chooses the point with the greatest EI to explore. After obtaining the optimal ratio $\boldsymbol{\lambda}^{*}$, we can determine the allocated rank: $r_{i}=(1-\lambda_{i})d_{1}d_{2}/(d_{1}+d_{2})$. To fully leverage the acceleration effect of GPU matrix multiplication, we adhere to Nvidia’s user guidelines [^2] https://docs.nvidia.com/deeplearning/performance/dl-performance-matrix-multiplication/index.html#gpu-imple by rounding the low-rank dimensions to the nearest multiple of eight.
The evaluation metric and validation data play a significant role in the optimization performance of BO. They must meet two criteria: cost-effectiveness and accurately reflect actual changes in performance. To this end, we propose a sensitive-based sampling method. This method randomly samples $n$ allocation schemes, calculates the variance of the perplexity of each sample under different allocations, and selects the top-$k$ samples as validation data. In addition, considering the smaller validation set may not comprehensively reflect the LLM’s performance, potentially leading to over-fitting in the validation set. To prevent BO from blindly improving the compressed model’s language modeling performance on the validation set, we aim to make the compressed model have a prediction distribution for the next word close to the original model. Hence, we employ the reverse KL divergence to quantify the difference:
@@ 4 Experiments
4.1 Baseline and Datasets
We compare our method with the competitive structured pruning and low-rank compression methods in LLMs: LLM-Pruner Ma et al. (2023), FLAP An et al. (2023), SliceGPT Ashkboos et al. (2024), LoRD Kaushal et al. (2023), ASVD Yuan et al. (2023). We provide the detailed description of baseline methods in Appendix B.
@@ 4.2 Experimental Details
In our main experiments, we apply our method to LLaMA-v2-7b and LLaMA-v2-13b. We randomly select 1,024 samples from the training set of C4 as the calibration data. Each sample has a sequence length of 4,096. To estimate the covariance matrix while saving memory usage, we employ the Welford’s online algorithm Welford (1962). For the pooled covariance matrix, we partition the calibrated data into 32 groups. During the Bayesian optimization, we utilize the Matern kernel as the covariance function.We randomly sample 20 low-rank allocation schemes and select the top-100 samples with greatest perplexity variance of Wikipedia as the evaluation data. Each sample has a sequence length of 4,096 (4k tokens). Considering that Bayesian optimization is not well-suited for high-dimensional scenarios, we conduct experiments with two settings based on the observations in Section 2.2: (a) $5\times 1$: We allow $attn\_q$ and $attn\_k$ to share a low-rank dimension, and the same type of parameters across different layers to also share a low-rank dimension, thus BO only optimizes 5 parameters; (b) $5\times 4$: Building on the setup of (a), we divide the model’s layers into 4 groups in sequence, with no parameter sharing between different groups, resulting in BO needing to optimize 20 parameters. Moreover, given that the parameters of the FFN module are more sensitive to low-rank compression than those of the attention module, we set the length scale for the attention and FFN parameters in the Matern kernel to 1.0 and 0.8, respectively, to emphasize the more significant impact of FFN parameters’ rank changes on model performance. We run 50 epochs BO to search the optimal low-rank allocation. At the post-training stage, following LLM-Pruner, we use the Alpaca dataset Taori et al. (2023) and train 2 epochs. More details can be found in the Appendix C.
@@ 4.3 Main Results
ASVD 73.61 71.93 66.05 64.17 65.24 36.26 37.40 59.24
Bolaco ($5\times 1$) 72.17 75.52 66.76 67.72 73.02 38.74 40.60 62.08
Bolaco ($5\times 1$ w/ PT) 73.79 77.53 72.72 68.11 73.19 42.24 43.60 64.45
Bolaco ($5\times 4$) 75.05 75.46 67.12 67.01 72.05 38.91 42.40 62.57
Bolaco ($5\times 4$ w/ PT) 75.84 76.61 71.70 65.67 72.60 41.81 45.00 64.18
@@
ASVD 59.42 55.93 35.05 52.25 34.30 26.45 26.60 41.43
Bolaco ($5\times 1$) 68.26 72.09 57.46 65.87 65.19 32.85 37.20 56.99
Bolaco ($5\times 1$ w/ PT) 70.34 74.32 67.81 65.04 69.02 38.31 41.80 60.95
Bolaco ($5\times 4$) 70.37 71.44 59.62 64.80 66.46 34.39 38.60 57.95
Bolaco ($5\times 4$ w/ PT) 71.83 75.19 68.03 65.67 69.15 38.74 42.40 61.57
Table 1: Zero-shot performance of the compressed LLaMA-v2-7b models. w/ PT means the method with post-training. Bold denotes the best result at the same compression ratio, while underline indicates the second best result.
We report the perplexity of language modeling for various compression methods at different compression ratios in Figure 3 and 6, and the zero-shot common sense reasoning results in Table 1 and 5 (in Appendix). In terms of language modeling capabilities, FLAP demonstrates strong competitiveness, particularly when the compression rate exceeds 30%, where FLAP’s perplexity is slightly better than our Bolaco ($5\times 1$). However, in the 7b model, Bolaco ($5\times 4$) achieves the best language modeling performance at high compression rates. Nevertheless, in the 13b model, despite Bolaco ($5\times 4$) still leading other compression techniques, it maintains a certain gap from FLAP. For zero-shot tasks, our method significantly outperforms all baselines without any further post-training, achieving an average performance increase of 1.5-2% across seven datasets. After post-training with only about 1‰ parameters and 3 hours, our method further narrows down the performance difference between the compressed model and the original model. It retains 96%-98% of the original model’s performance at the 20% compression ratio, and at a 30% compression ratio, it maintains 91%-95% of the performance. Comparing the $5\times 1$ and $5\times 4$ setting, we find that the performance difference between the two is not significant. At the 20% compression ratio, simply allocating different low-rank dimensions to different types of parameters suffices to achieve the best current performance. However, at the 30% compression rate, the $5\times 4$ setting outperforms the $5\times 1$, indicating that more granular low-rank assignments contribute to enhanced performance in compressed models at higher compression rates.
@@ 5.2 Impact of Objective Function
PPL ($5\times 1$) 8.36 48.42 61.70
w/ RKL ($5\times 1$) 8.27 47.06 62.08
PPL ($5\times 4$) 8.07 47.96 60.98
w/ RKL ($5\times 4$) 7.96 45.84 62.57
@@
PPL ($5\times 1$) 13.78 71.50 56.97
w/ RKL ($5\times 1$) 13.41 70.52 56.99
PPL ($5\times 4$) 12.65 68.85 57.57
w/ RKL ($5\times 4$) 13.70 72.14 57.95
@@
We explore the impact of the objective function in the Bayesian optimization stage. We conduct experiments on LLaMA-v2-7b and report results in Table 3. Overall, incorporating the reverse KL divergence (RKL) between the compressed model and the original model’s predictive distribution into the objective function can lead to a better low-rank dimensions allocation. Especially in the $5\times 4$ setting, which is more difficult to optimize for Bayesian optimization, the performance gains from RKL term are even more obvious. We suppose that the RKL term may serve two roles. Firstly, as a regularization term, it prevents overfitting on smaller validation sets during BO. Although the compressed model exhibits a slight increase in perplexity on the language modeling dataset at the 20% compression rate with the $5\times 4$ setting, there is a significant improvement in performance on downstream tasks. Secondly, incorporating the RKL term may smooth the objective function, enabling the Gaussian process surrogate model to more accurately approximate the real black-box objective function.